OpenStudy (anonymous):

The leg of a right triangle is 5 units and the hypotenuse is 8 units. What is the length, in units, of the other leg of the triangle? units 39 units units 89 units

OpenStudy (kobeni-chan):

To find the lengths of a triangle, set it up like this: a^2 + b^2 = c^2. a and b are the legs of a triangle, and c is the hypotenuse 5^2 + b^2 = 8^2

OpenStudy (kobeni-chan):

you'll then have 25 + b^2 = 64. You can change this to 64 - 25 = b^2. Find what the difference of 64 and 25 is, then square root it.

OpenStudy (anonymous):

i get 39 and squar root is 6.2 ????

OpenStudy (kobeni-chan):

yup, as long as that's an option

OpenStudy (anonymous):

yep it is thanks

OpenStudy (kobeni-chan):

your welcome :)